# Notions of entropy for unitary representations

### v2

@@ 1.4 Almost periodic entropy
Since its introduction in [Bow10b], Bowen’s notion of sofic entropy has taken a central place in the ergodic theory of actions of non-amenable groups. This is the final notion of entropy that we pursue across the analogy between ergodic theory and representation theory in this paper.
Consider again a countable group $\Gamma$. It is sofic if it admits a certain kind of approximation by ‘near actions’ by permutations of finite sets. This property is introduced by Gromov in [Gro99], and first called ‘sofic’ by Weiss in [Wei00]. Given a finite set $A$ and a stationary $A$-valued process over $\Gamma$, Bowen’s idea is to fix such a sequence of approximating near actions on finite sets $V_{1}$, $V_{2}$, …, and then count the $A$-colourings of each $V_{n}$ whose ‘local statistics’ under those near actions are close to the law of the process. Very roughly, the lim-sup exponential growth rate of these counts is the definition of the sofic entropy of the process along that sequence of near actions. The surveys [Bow18, Bow20] offer a thorough account.
The next innovation in the present paper is an analogous construction for a unitary representation $\pi$ and a tuple $v_{1},\dots,v_{k}\in H_{\pi}$. This time we approximate by a sequence of finite-dimensional representations, and for each of these we consider the vectors that give roughly the same positive definite function on $\Gamma$ as $v_{1}$, …, $v_{k}$ do in $\pi$.
In pursuing this idea, we quickly find that it is easier to allow greater generality still. We fix a separable, unital C*-algebra $\mathfrak{A}$ and consider all its representations on separable Hilbert spaces. The case of a group $\Gamma$ is recovering by taking $\mathfrak{A}$ equal to $C^{\ast}\Gamma$. However, even in that case certain helpful auxiliary constructions lead us to consider other C*-algebras as well.
To explain our definition more precisely, consider a representation $\pi$ of $\mathfrak{A}$ and vectors $v_{1}$, …, $v_{k}\in H_{\pi}$. Define their type to be the $\mathrm{M}_{k}$-valued map
@@
Imagining that $O$ is a small neighbourhood of a particular completely positive map $\varphi$, this is the analog of a set of ‘good models’ for a given shift-invariant measure in sofic entropy theory.
Finally, consider a sequence $\boldsymbol{\pi}:=(\pi_{n})_{n\geq 1}$ of representations of $\mathfrak{A}$ whose dimensions $d_{n}$ are finite but diverge. We refer to it as an almost periodic sequence for $\mathfrak{A}$. See Definition 8.17 and the rest of Subsection 8.4, which also discusses possible modes of convergence for such a sequence.
@@
where $O$ ranges over neighbourhoods of $\varphi$, $\mathrm{vol}_{2kd_{i}}$ refers to Lebesgue measure in $2kd_{i}$ real dimensions, and $v(d_{i})$ is the volume of the unit ball in $\mathbf{C}^{d_{i}}$. A full explanation is given in Definition 9.4. The formula above is a direct analog of the usual definition of sofic entropy for a measure-preserving system with a distinguished generating observable.
@@
Our final main result is a formula for $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)$ as a Fuglede–Kadison determinant. It holds whenever $\varphi$ is ‘asymptotically associated’ to $\boldsymbol{\pi}$ and the pulled-back traces $d_{i}^{-1}\mathrm{tr}_{d_{i}}\circ\pi_{i}$ converge to a limiting tracial state $\tau$ of $\mathfrak{A}$. ‘Asymptotic association’ means that, for every neighbourhood $O$ of $\varphi$, the set $\mathcal{X}(\pi_{i},O)$ is nonempty for infinitely many $i$; without this assumption $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)$ is simply forced to be $\infty$- ∞.
@@
Let us emphasize two related features of Theorem C that are substantially different from Theorems A and B:
@@
  • Theorem C does not make any assumption on $\mathfrak{A}$ itself that corresponds to soficity of a group. However, this assumption is implicit in the hypothesis that $\tau$ is a limit of normalized finite-dimensional traces. By applying this theorem to $C^{\ast}\Gamma$ when $\Gamma$ is a free group and lifting $\tau$ from a quotient group of $\Gamma$, one recovers a theorem for precisely the class of sofic groups: see Subsection 8.4.
In comparison with Theorem S, Theorem C has the interesting new feature that $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)$ may equal $\infty$- ∞ if $\varphi$ is not asymptotically associated to $\boldsymbol{\pi}$, even though $\log\Delta_{\tau}\varphi_{\mathrm{ac}}$ may still be finite in that case. If $\mathfrak{A}=C^{\ast}\Gamma$ and $\tau$ is the state given by the regular character of $\Gamma$, then this is possible only if $\Gamma$ is non-amenable, and reflects basic features of the representation theory of non-amenable groups.
@@
As far as I know, the nearest precursors to Theorem C in the literature are formulas for the sofic entropy of certain special measure-preserving systems in [Lyo05, Lyo10] and especially [Hay16, Hay21b]. We compare our work with these in Subsection 9.7.
@@ 2 Notation, conventions, and basic facts
This chapter sets some notation that we use throughout, some of which is slightly non-standard. Apart from that, most readers should be able to skip it and then refer back to it as necessary.
@@ 2.1 Linear algebra
We assume standard ideas from linear algebra and matrix analysis. For definiteness I use [HJ13] as a reference wherever possible
@@
At some points below it is important to distinguish between row and column vectors. In those cases we regard the elements of $\mathbf{C}^{k}$ as row vectors by default, and write $\mathbf{C}^{\oplus k}$ for the space of height-$k$ column vectors when these are needed specifically. More generally, if $H$ is a vector space, then we write $H^{\oplus k}$ for the $k$-fold inflation of $H$, which is the vector space of height-$k$ column vectors with entries in $H$. If $S$ is a set, possibly infinite, and $H$ is a Hilbert space, then we extend this notation further by writing $H^{\oplus S}$ for the Hilbert-space direct sum of an $S$-indexed family of copies of $H$, still regarded as a space of column vectors. This insistence on column vectors is slightly unusual in functional analysis, but for finite $k$ it enables us to use matrix-vector notation from linear algebra in places where it simplifies the exposition.
We write $\mathrm{M}_{n,k}$ for the space of $n$-by-$k$ matrices over the complex numbers, and identify these with linear maps from $\mathbf{C}^{\oplus k}$ to $\mathbf{C}^{\oplus n}$ using matrix-vector multiplication. By writing such a matrix as $[v_{1},\dots,v_{k}]$, where $v_{1}$, …, $v_{k}$ are its columns, we identify it with a $k$-tuple of vectors in $\mathbf{C}^{\oplus n}$. We generalize this notation further by allowing columns from a general vector space $H$, meaning that a linear map from $\mathbf{C}^{\oplus k}$ to $H$ may still be written in the form $[v_{1},\dots,v_{k}]$. Moreover, we sometimes abuse notation by calling the linear map $V$ itself a ‘$k$-tuple of vectors in $H$’. If $H$ is an inner product space, the adjoint $V^{\ast}$ is the map from $H$ to $\mathbf{C}^{\oplus k}$ whose output coordinates are given by the inner products with the vectors $v_{i}$.
@@
We write $\mathrm{M}_{k+}$ for the closed cone of positive semidefinite members of $\mathrm{M}_{k}$. It defines the positive definite ordering $\leq_{\mathrm{pd}}$ on self-adjoint matrices (also called the ‘Loewner ordering’). If $Q\in\mathrm{M}_{K+}$, then its determinant and trace are related by the inequality
@@
This is the inequality of arithmetic and geometric means applied to the eigenvalues of $Q$.
@@
If $P$ is an orthogonal projection in a Hilbert space $H$, then we sometime write $P^{\perp}$ for $I-P$.
@@ 2.2 Gram matrices and Schur complements
for some $k$-tuple $Y$ in $M$, so $PY=0$. The Gram matrix of this perturbation is
@@
In particular, this shows that $PV$ uniquely minimizes the Gram matrix $Z^{\ast}Z$ among all such perturbation in the positive definite ordering. It therefore also has minimal determinant among the Gram matrices of these perturbations [HJ13, Corollary 7.7.4(e)]. The Gram matrix of $PV$ is an abstract form of the Schur complement of $V^{\ast}V$ relative to $M$. If $M$ is the span of another tuple $U=[u_{1},\dots,u_{\ell}]$, and we write $W$ for the combined tuple $[u_{1},\dots,u_{\ell},v_{1},\dots,v_{k}]$, then the Schur complement participates in Schur’s determinantal formula:
@@ 2.4 Measure and integration in high dimensions
For any dimension $d$, we write $\mathrm{vol}_{d}$ for Lebesgue measure on $\mathbf{R}^{d}$. For any positive integers $d$ and $k$, we also write $\mathrm{vol}_{2kd}$ for the measure on $\mathrm{M}_{k,d}(\mathbf{C})$ obtained by identifying this space with $\mathbf{R}^{2kd}$. We write $\mathrm{S}^{d-1}$ for the unit sphere in $\mathbf{R}^{d}$ or in any other vector space that has a standing identification with $\mathbf{R}^{d}$. We write $\sigma_{d-1}$ for the surface-area measure on $\mathrm{S}^{d-1}$ normalized to have total mass $1$, and we refer to an integral with respect to $\sigma_{d-1}$ as a ‘spherical average’.
@@ 3 Operator algebras, C*-algebras and representations
Although C*-algebras are the natural generality for much of our work below, I have tried to avoid assuming very advanced parts of that theory. This section recalls most of the background we need. Where possible I cite the standard texts [Con90, RS72] on functional analysis and [Con00, Dix77, Dix81] on C*-algebras and operator algebras. I cite various specific results from original papers.
Readers familiar with these topics may wish to skip this preliminary chapter and refer back to it as necessary.
@@ 3.1 C*-algebras, von Neumann algebras, and affiliated operators
Our guiding example is the group C*-algebra $C^{\ast}\Gamma$ of a countable group $\Gamma$. This is the maximal C*-completion of the group algebra $\mathbf{C}[\Gamma]$, so it is indeed separable and unital. We recall some more specialized results about group C*-algebras in Section 5 below. Our results about almost periodic entropy in Section 9 could also apply to the unital augmentation of the group C*-algebra of a locally compact, second countable group.
@@
Some of the constructions we need later involve operators on a Hilbert space that are closed and densely defined but unbounded. The basic theory of these can be found in [Con90, Chapter X], including the unbounded version of the spectral theorem [Con90, Section X.4] which leads to an unbounded extension of the polar decomposition [Con90, Exercise X.4.6] (or see [RS72, Section VIII.9]).
A closed and densely defined operator $T$ is affiliated to the von Neumann algebra $\mathfrak{N}$ if it commutes with every unitary element of $\mathfrak{N}^{\prime}$: see [Dix81, Exercise I.1.10]. This includes the assertion that those unitary elements preserve $\mathrm{dom}\,T$. The collection of operators affiliated to $\mathfrak{N}$ is closed under forming adjoints, and if such an operator $T$ has polar decomposition $U|T|$, then the uniqueness of that decomposition implies that $U\in\mathfrak{N}$ and that $|T|$ is also affiliated to $\mathfrak{N}$. If $\mathfrak{N}$ is a finite von Neumann algebra, then sums and products of affiliated operators are also still affiliated operators, provided each of these constructions is following by taking graph closures [Dix81, Exercise III.1.13.c].
@@ 3.2 Representations and positive functionals
We use $\oplus$ to denote direct sums in the categories of Hilbert spaces or representations [Dix77, Subsection 2.2.3], and we use $\otimes$ for tensor products of Hilbert spaces or of operators or von Neumann algebras on them [Dix81, Sections I.2.3–4]. We do not need the more involved theory of tensor products of abstract C*-algebras. For a representation $\pi$ and any positive integer $k$, we write either $\pi^{\oplus k}$ or $\pi\otimes I_{k}$ for the direct sum of $k$ copies of $\pi$. We extend this notation to allow $k=\infty$ by taking $I_{\infty}$ to be the identity on some fixed choice of infinite-dimensional separable complex Hilbert space. In either case we call $\pi^{\oplus k}$ the $k$-fold inflation of $\pi$ (this construction is called ‘ampliation’ in [Dix77]).
@@
  • $\pi$ is equivalent to $\rho$, written $\pi\cong\rho$, if there is a unitary operator from $H_{\pi}$ to $H_{\rho}$ that intertwines $\pi$ with $\rho$;
@@
Another possible relation is that $\pi$ is contained in an inflation of $\rho$, which we can always write as $\pi\lesssim\rho^{\oplus\infty}$.
@@
Now let $\kappa:=\pi\oplus\rho$. Regard $H_{\pi}$ and $H_{\rho}$ as invariant subspaces of $H_{\kappa}$, and let $P$ be the orthogonal projection from $H_{\kappa}$ onto $H_{\pi}$. The next result is [Dix77, Proposition 5.2.4].
@@
defines a positive linear functional on $\mathfrak{A}$. All positive linear functionals arise this way because of the GNS construction [Dix77, Section 2.4]. We denote the GNS representation of $\varphi$ by $\pi_{\varphi}$. We write that the functional in (3.1) is associated to $\pi$ by $v$. By the uniqueness of the GNS construction, $\varphi$ is associated to $\pi$ if and only if $\pi_{\varphi}$ is contained in $\pi$. Alternatively, if the representation $\pi$ is understood, we adapt a term from information theory by calling $\Phi^{\pi}_{v}$ the type of $v$ (compare [CT06, Section 11.1], for example).
We write $\mathfrak{A}^{\ast}_{+}$ for the space of positive linear functionals on $\mathfrak{A}$, and always endow it with the weak topology inherited from $\mathfrak{A}$. We write $\Sigma(\mathfrak{A})$ for the further subset of positive functionals $\varphi$ that satisfy $\varphi(1)=1$. Those functionals are called states, and $\Sigma(\mathfrak{A})$ is the state space of $\mathfrak{A}$. These standard notions and their relation to representations can be found in [Dix77, Chapter 2].
Given any two continuous linear functionals $\varphi$ and $\psi$ on $\mathfrak{A}$, we write $\varphi\leq\psi$ if the functional $\psi-\varphi$ is positive. This defines a partial order on $\mathfrak{A}^{\ast}$ in which $\mathfrak{A}^{\ast}_{+}$ is the non-negative cone. This relation implies containment of GNS representations by the following basic result (see [Dix77, Proposition 2.5.1]).
@@
We need two more relations involving positive functionals and representations:
  • If $\pi$ is a representation and $\varphi$ is a positive functional, then $\varphi$ is $\pi$-normal if $\varphi$ is associated to $\pi^{\oplus\infty}$. This is equivalent to $\varphi$ having the form $\widetilde{\varphi}\circ\pi$ for some normal positive functional $\widetilde{\varphi}$ on $\pi(\mathfrak{A})^{\prime\prime}$.
@@
Proposition 3.3 is from [Tak58]. The heart of the proof is a corresponding decomposition of $\pi_{\varphi}$ relative to $\pi$. That decomposition can also be found as [Mac76, Theorem 1.11] (phrased there in the special case of group representations, but with a proof that works in general).
If $X$ is a compact metrizable space and $\mathfrak{A}=C(X)$, then the Riesz–Markov theorem identifies positive linear functionals on $\mathfrak{A}$ with finite Borel measures on $X$. With this in hand, routine steps reduce Proposition 3.3 to the Lebesgue decomposition from measure theory. With this in mind, we call (3.2) the Lebesgue decomposition of $\varphi$ relative to $\pi$.
Starting from that example, the study of functionals on a C*-algebra or von Neumann algebra has long been regarded as a ‘non-commutative’ extension of the theory of measure and integration. See [Seg53b, Seg53a] and [Dix53] for early developments of this point of view.
To accompany the Lebesgue decomposition, there are also non-commutative extensions of the Radon–Nikodym theorem. These describe the structure of $\pi$-normal functionals in terms of vectors or operators associated to $\pi$. There are actually several such theorems that allow for the non-commutativity of $\mathfrak{A}$ in different ways. Proposition 3.2 is sometimes classed as one of these theorems, although the assumption $\varphi\leq c\psi$ is typically much stronger than $\varphi$ being $\pi_{\psi}$-normal. In this work we also need another such theorem, but it depends on the extra structure of a tracial vector: see Proposition 3.5 below.

### v3

@@ 1.4 Almost periodic entropy
Since its introduction in [25], Bowen’s notion of sofic entropy has taken a central place in the ergodic theory of actions of non-amenable groups. This is the final notion of entropy that we pursue across the analogy between ergodic theory and representation theory in this paper.
Consider again a countable group $\Gamma$. It is sofic if it admits a certain kind of approximation using ‘near actions’ by permutations of finite sets. This property was introduced by Gromov in [76], and first called ‘sofic’ by Weiss in [159]. Given a finite set $A$ and a stationary $A$-valued process over $\Gamma$, Bowen’s idea was to fix such a sequence of approximating near actions on finite sets $V_{1}$, $V_{2}$, …, and then count the $A$-colourings of each $V_{n}$ whose ‘local statistics’ under those near actions are close to the law of the process. Very roughly, the lim-sup exponential growth rate of these counts defines the sofic entropy of the process along that sequence of near actions. The surveys [28, 26] offer a thorough account.
Our next subject is an analogous construction for a unitary representation $\pi$ of $\Gamma$ and a tuple $v_{1},\dots,v_{k}\in H_{\pi}$. This time we approximate by a sequence of finite-dimensional representations, and for each of these we consider the vectors that give roughly the same positive definite function on $\Gamma$ as $v_{1}$, …, $v_{k}$ do in $\pi$.
In pursuing this idea, we quickly find that it is easier to allow greater generality still. We fix a separable, unital C*-algebra $\mathfrak{A}$ and consider all its unital representations on separable Hilbert spaces. The case of a group $\Gamma$ is recovering by taking $\mathfrak{A}$ equal to $C^{\ast}\Gamma$. However, even in that case certain helpful auxiliary constructions lead us to consider other C*-algebras as well.
Given a representation $\pi$ of $\mathfrak{A}$ and vectors $v_{1}$, …, $v_{k}\in H_{\pi}$, define their type to be the $\mathrm{M}_{k}$-valued map
@@
Imagining that $O$ is a small neighbourhood of a particular map $\varphi$, this is the analog of a set of ‘good models’ for a given shift-invariant measure in sofic entropy theory.
Finally, consider a sequence $\boldsymbol{\pi}=(\pi_{n})_{n\geq 1}$ of representations of $\mathfrak{A}$ whose dimensions $d_{n}$ are finite but diverge. We refer to it as an almost periodic sequence for $\mathfrak{A}$. See Subsection 8.4, which also discusses possible modes of convergence for such a sequence.
@@
where $O$ ranges over neighbourhoods of $\varphi$, $\mathrm{vol}_{2kd_{i}}$ refers to Lebesgue measure in $2kd_{i}$ real dimensions, and $v(d_{i})$ is the volume of the unit ball in $\mathbf{C}^{d_{i}}$. See Definition 9.4 for a full explanation. The formula above is a direct analog of the usual definition of sofic entropy for a finite-valued stationary process over a countable sofic group.
@@
Our final main result is a formula for $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)$ as a Fuglede–Kadison determinant. It holds whenever $\varphi$ is ‘asymptotically associated’ to $\boldsymbol{\pi}$ and the pulled-back traces $d_{i}^{-1}\mathrm{tr}_{d_{i}}\circ\pi_{i}$ converge to a limiting tracial state $\tau$ of $\mathfrak{A}$. ‘Asymptotic association’ means that, for every neighbourhood $O$ of $\varphi$, the set $\mathcal{X}(\pi_{i},O)$ is nonempty for infinitely many $i$; if this fails then $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)$ is simply forced to be $\infty$- ∞.
@@
Let us emphasize two features of Theorem C that are substantially different from Theorems A and B:
@@
  • Theorem C does not make any assumption on $\mathfrak{A}$ itself that corresponds to soficity of a group. However, such an assumption is implicit in the hypothesis that $\tau$ is a limit of normalized finite-dimensional traces. By applying this theorem to $C^{\ast}\Gamma$ when $\Gamma$ is a free group and $\tau$ is lifted from a quotient group of $\Gamma$, one recovers a theorem for precisely Rădulescu’s class of hyperlinear groups from [125]: see Subsection 8.4.
In comparison with Theorem S, Theorem C has the interesting new feature that $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)$ may equal $\infty$- ∞ if $\varphi$ is not asymptotically associated to $\boldsymbol{\pi}$, even though $\log\Delta\varphi_{\mathrm{ac}}$ may still be finite in that case. If $\mathfrak{A}=C^{\ast}\Gamma$ and $\tau$ is the state given by the regular character of $\Gamma$, then this is possible only if $\Gamma$ is non-amenable, and reflects basic features of the representation theory of non-amenable groups.
@@
As far as I know, the nearest precursors to Theorem C in the literature are formulas for the sofic entropy of certain special measure-preserving systems in [110, 111] and especially [80, 84]. We compare our work with these in Subsection 9.7.
@@ 2 Notation, conventions, and basic facts
This section sets some notation that we use throughout, some of which is slightly non-standard. Readers may prefer to skip it and then refer back to it as necessary.
@@ 2.1 Linear algebra
We assume standard ideas from linear algebra and matrix analysis. For definiteness I use [88] as a reference wherever possible
@@
At some points below it is convenient to distinguish between row and column vectors. In those cases we regard the elements of $\mathbf{C}^{k}$ as row vectors by default, and write $\mathbf{C}^{\oplus k}$ for the space of height-$k$ column vectors when these are needed specifically. More generally, if $H$ is a vector space, then we write $H^{\oplus k}$ for the $k$-fold inflation of $H$, which is the vector space of height-$k$ column vectors with entries in $H$. If $S$ is a set, possibly infinite, and $H$ is a Hilbert space, then we extend this notation further by writing $H^{\oplus S}$ for the Hilbert-space direct sum of an $S$-indexed family of copies of $H$, still regarded as a space of column vectors. This insistence on column vectors is slightly unusual in functional analysis, but for finite $k$ it enables us to use matrix-vector notation from linear algebra in places where it simplifies the exposition.
We write $\mathrm{M}_{n,k}$ for the space of $n$-by-$k$ matrices over the complex numbers, and identify these with linear maps from $\mathbf{C}^{\oplus k}$ to $\mathbf{C}^{\oplus n}$ using matrix-vector multiplication. By writing such a matrix as $[v_{1},\dots,v_{k}]$, where $v_{1}$, …, $v_{k}$ are its columns, we can identify it with a $k$-tuple of vectors in $\mathbf{C}^{\oplus n}$. We generalize this notation further by allowing columns from any vector space $H$, so a linear map from $\mathbf{C}^{\oplus k}$ to $H$ may still be written in the form $[v_{1},\dots,v_{k}]$. Moreover, we sometimes abuse notation by calling the linear map $V$ itself a ‘$k$-tuple of vectors in $H$’. If $H$ is an inner product space, then the adjoint $V^{\ast}$ is the map from $H$ to $\mathbf{C}^{\oplus k}$ whose output coordinates are given by the inner products with the vectors $v_{i}$.
@@
We write $\mathrm{M}_{k+}$ for the closed cone of positive semidefinite members of $\mathrm{M}_{k}$. It defines the positive definite ordering on self-adjoint matrices (also called the ‘Loewner ordering’). If $Q\in\mathrm{M}_{k+}$, then its determinant and trace are related by the inequality
@@
This is simply the inequality of arithmetic and geometric means applied to the eigenvalues of $Q$.
@@
If $P$ is an orthogonal projection in a Hilbert space $H$, then we use $P^{\perp}$ as a shorthand for $I-P$.
@@ 2.2 Gram matrices and Schur complements
for some $k$-tuple $Y$ in $M$. The Gram matrix of this perturbation is
@@
In particular, this shows that $PV$ uniquely minimizes the Gram matrix $Z^{\ast}Z$ among all such perturbation in the positive definite ordering. It therefore also has minimal determinant among the Gram matrices of these perturbations [88, Corollary 7.7.4(e)]. The Gram matrix of $PV$ is an abstract form of the Schur complement of $V^{\ast}V$ relative to $M$. If $M$ is the span of another tuple $U=[u_{1},\dots,u_{\ell}]$, and we write $W$ for the combined tuple $[u_{1},\dots,u_{\ell},v_{1},\dots,v_{k}]$, then the Schur complement participates in Schur’s determinantal formula:
@@ 2.4 Measure and integration in high dimensions
For any dimension $d$, we write $\mathrm{vol}_{d}$ for Lebesgue measure on $\mathbf{R}^{d}$. For any positive integers $d$ and $k$, we also write $\mathrm{vol}_{2kd}$ for the measure on $\mathrm{M}_{k,d}$ obtained by identifying this space with $\mathbf{R}^{2kd}$. We write $\mathrm{S}^{d-1}$ for the unit sphere in $\mathbf{R}^{d}$ or in any other vector space that has a standing identification with $\mathbf{R}^{d}$. We write $\sigma_{d-1}$ for the surface-area measure on $\mathrm{S}^{d-1}$ normalized to have total mass $1$, and we refer to an integral with respect to $\sigma_{d-1}$ as a ‘spherical average’.
@@ 3 Operator algebras, C*-algebras and representations
Although C*-algebras are the natural generality for much of our work below, I have tried to avoid assuming very advanced parts of that theory. This section recalls most of the background we need. Where possible I cite the standard texts [37, 124] on functional analysis and [38, 48, 49] on C*-algebras and operator algebras. I cite a few specific results from original papers.
Readers familiar with these topics may wish to skip this section and refer back to it as necessary.
@@ 3.1 C*-algebras, von Neumann algebras, and affiliated operators
Our guiding example is the group C*-algebra $C^{\ast}\Gamma$ of a countable group $\Gamma$. This is the maximal C*-completion of the group algebra $\mathbf{C}[\Gamma]$, so it is indeed separable and unital. Our results about almost periodic entropy in Section 9 could also apply to the unital augmentation of the group C*-algebra of a locally compact, second countable group.
@@
Some of the constructions we need later involve operators on a Hilbert space that are closed and densely defined but unbounded. The basic theory of these can be found in [37, Chapter X], including the unbounded version of the spectral theorem [37, Section X.4] which leads to an unbounded extension of the polar decomposition [37, Exercise X.4.6] (or see [124, Section VIII.9]).
A closed and densely defined operator $T$ is affiliated to the von Neumann algebra $\mathfrak{N}$ if it commutes with every unitary element of $\mathfrak{N}^{\prime}$: see [49, Exercise I.1.10]. This includes the assertion that those unitary elements preserve $\mathrm{dom}\,T$. The collection of operators affiliated to $\mathfrak{N}$ is closed under forming adjoints, and if such an operator $T$ has polar decomposition $U|T|$, then the uniqueness of that decomposition implies that $U\in\mathfrak{N}$ and that $|T|$ is also affiliated to $\mathfrak{N}$. If $\mathfrak{N}$ is a finite von Neumann algebra, then sums and products of affiliated operators are also still affiliated operators, provided each of these constructions is followed by taking graph closures [49, Exercise III.1.13.c].
@@ 3.2 Representations and positive functionals
We use $\oplus$ to denote direct sums in the categories of Hilbert spaces or representations [48, Subsection 2.2.3], and we use $\otimes$ for tensor products of Hilbert spaces or of operators or von Neumann algebras on them [49, Sections I.2.3–4]. We do not need the more involved theory of tensor products of abstract C*-algebras. For a representation $\pi$ and any positive integer $k$, we write either $\pi^{\oplus k}$ or $\pi\otimes I_{k}$ for the direct sum of $k$ copies of $\pi$. We extend this notation to allow $k=\infty$ by taking $I_{\infty}$ to be the identity on some fixed choice of infinite-dimensional separable complex Hilbert space. We refer to $\pi^{\oplus k}$ as the $k$-fold inflation of $\pi$ (this construction is called ‘ampliation’ in [48]).
@@
  • $\pi$ is equivalent to $\rho$, written $\pi\simeq\rho$, if there is a unitary operator from $H_{\pi}$ to $H_{\rho}$ that intertwines $\pi$ with $\rho$;
@@
Another possible relation is that $\pi$ is contained in an inflation of $\rho$, which we can always write as $\pi\lesssim\rho^{\oplus\infty}$ if both representations are separable.
@@
Now let $\kappa:=\pi\oplus\rho$. Regard $H_{\pi}$ and $H_{\rho}$ as invariant subspaces of $H_{\kappa}$, and let $P$ be the orthogonal projection from $H_{\kappa}$ onto $H_{\pi}$. The next result is [48, Proposition 5.2.4].
@@
defines a positive linear functional on $\mathfrak{A}$. We write that the functional in (3.1) is associated to $\pi$ by $v$. Alternatively, if the representation $\pi$ is understood, we adapt a term from information theory by calling $\Phi^{\pi}_{v}$ the type of $v$ (compare [40, Section 11.1], for example).
All positive linear functionals on $\mathfrak{A}$ arise this way because of the GNS construction [48, Section 2.4]. We denote the GNS representation of $\varphi$ by $\pi_{\varphi}$. By the uniqueness of the GNS construction, $\varphi$ is associated to $\pi$ if and only if $\pi_{\varphi}$ is contained in $\pi$.
We write $\mathfrak{A}^{\ast}_{+}$ for the space of positive linear functionals on $\mathfrak{A}$, and always endow it with the weak topology inherited from $\mathfrak{A}$. We write $\Sigma(\mathfrak{A})$ for the further subset of positive functionals $\varphi$ that satisfy $\varphi(1)=1$. Those functionals are called states, and $\Sigma(\mathfrak{A})$ is the state space of $\mathfrak{A}$. These standard notions and their relation to representations can be found in [48, Chapter 2].
Given any two continuous linear functionals $\varphi$ and $\psi$ on $\mathfrak{A}$, we write $\varphi\leq\psi$ if the functional $\psi-\varphi$ is positive. This defines a partial order on $\mathfrak{A}^{\ast}$ in which $\mathfrak{A}^{\ast}_{+}$ is the non-negative cone. For positive functionals, this relation implies containment of GNS representations by the following basic result (see [48, Proposition 2.5.1]).
@@
We also need two other relations involving positive functionals and representations:
  • If $\pi$ is a representation and $\varphi$ is a positive functional, then $\varphi$ is $\pi$-normal if $\varphi$ is associated to $\pi^{\oplus\infty}$. This is equivalent to $\varphi$ having the form $\widetilde{\varphi}\circ\pi$ for some normal positive functional $\widetilde{\varphi}$ on $\pi(\mathfrak{A})^{\prime\prime}$, by the general characterization of normal positive functionals on a von Neumann algebra [38, Theorem 46.4].
@@
Proposition 3.3 is from [141]. The heart of the proof is a corresponding decomposition of $\pi_{\varphi}$ relative to $\pi$. That decomposition can also be found as [112, Theorem 1.11] (phrased there in the special case of group representations, but with a proof that works in general).
If $X$ is a compact metrizable space and $\mathfrak{A}=C(X)$, then the Riesz–Markov theorem identifies positive linear functionals on $\mathfrak{A}$ with finite Borel measures on $X$. Through this identification, routine steps reduce Proposition 3.3 to the Lebesgue decomposition from measure theory. With this in mind, we call (3.2) the Lebesgue decomposition of $\varphi$ relative to $\pi$. This also motivates the choice of subscripts for the two components in (3.2).
Starting from that example, the study of functionals on a C*-algebra or von Neumann algebra has long been regarded as a ‘non-commutative’ extension of the theory of measure and integration. See [132, 131] and [47] for early developments of this point of view.
To accompany the Lebesgue decomposition, there are also non-commutative extensions of the Radon–Nikodym theorem. These describe the structure of $\pi$-normal functionals in terms of vectors or operators related to $\pi$. There are actually several such theorems that allow for the non-commutativity of $\mathfrak{A}$ in different ways. Proposition 3.2 is sometimes classed as one of these theorems, although the assumption $\varphi\leq c\psi$ is typically much stronger than $\varphi$ being $\pi_{\psi}$-normal. In this work we also need another such theorem, but it depends on the extra structure of a tracial vector: see Proposition 3.5 below.